Anderson Family Steakhouse offers a variety of low-cost meals and quick service. Other than management, the Steakhouse operates with full time employees work 8 hrs/ day. The rest of employees are part time employees who are scheduled for 4 hrs shifts during peak meal times. On the Saturdays the Steakhouse is open from 11:00AM to10:00PM management want to develop a schedule for part time employees that will minimize labor cost and provide excellent customer service. The average wage rate for the part time employees is $7.60/ hr. The total number of full time and part-time employees needed varies with time of a day as shown.

Time

Total number of employees needed

11:00AM-Noon

9

Noon- 1:00PM

9

1:00PM-2:00PM

9

2:00PM-3:00PM

3

3:00PM-4:00PM

3

4:00PM-5:00PM

3

5:00PM-6:00PM

6

6:00PM-7:00PM

12

7:00PM-8:00PM

12

8:00PM-9:00PM

7

9:00PM-10:00PM

7

One full time employee comes on duty at 11:00 AM works 3 hrs.takes one hr. off, and returns for another 5 hrs. The other full-time employee comes to work at 1:00 PM and works the same 3 hours-on, 1-hour-off, 5-hours-on pattern.

  1. Write the complete linear programming model(objective function and constraints) and develop a minimum cost schedule for part time employees?

  1. What is total payroll for the part time employees? How many part time shifts are needed? (Use the surplus variables to comment on the desirability of scheduling at least some of the part time employees for 3 hrs shifts.)

  1. The restaurant wants to open an extra hour from 10 PM-11 PM with six part time employees needed during this hour. How much does payroll cost change?
  1. Assume that part-time employees can be assigned either 3hr or 4 hr shifts. Write the complete linear programming model and develop a minimum cost schedule for the part time employees. How many part time shifts are needed and what is the cost saving compared to the previous schedule? (restaurant closing time is 10 PM)

  1. The restaurant decides to give part-time employees a raise to $8.50 per hour. How much does the raise cost in part d?
Answer #1

a) Linear programming model is following:

Let

X1 = number of part-time employees coming to duty at 11 am

X2 = number of part-time employees coming to duty at 12 pm

X3 = number of part-time employees coming to duty at 1 pm

X4 = number of part-time employees coming to duty at 2 pm

X5 = number of part-time employees coming to duty at 3 pm

X6 = number of part-time employees coming to duty at 4 pm

X7 = number of part-time employees coming to duty at 5 pm

X8 = number of part-time employees coming to duty at 6 pm

Objective: Min 7.6*4*(X1+X2+X3+X4+X5+X6+X7+X8)

s.t.

X1 >= 9-1

X1+X2 >= 9-1

X1+X2+X3 >= 9-2

X1+X2+X3+X4 >= 3-1

X2+X3+X4+X5 >= 3-2

X3+X4+X5+X6 >= 3-1

X4+X5+X6+X7 >= 6-2

X5+X6+X7+X8 >= 12-2

X6+X7+X8 >= 12-2

X7+X8 >= 7-1

X8 >= 7-1

X1, X2, X3, X4, X5, X6, X7, X8 >= 0

b) Solution of the Linear program using LINGO is following:

Lingo 17.0 Solution Report Lingol File Edit Solver Window Help Lindo Model Lingol 3 回Solution Report . Lingo! Global optimal

Variable X1 X2 X3 value 8.000000 1.000000 0.000000 0.000000 0.000000 4.000000 0.000000 6.000000 Reduced Cost 0.000000 0.00000

Total payroll for the part-time employees = $ 577.6

Surplus variables for shift 2, 3, 4, 6 are 1,2,7,2 respectively.

Therefore, these many number of employees can be scheduled for 3 hrs shift.

c) The revised LP and its solution is following:

Lingo 170-Solution Report . Lingo! File Edit Solver Window Help Lindo Model-Lingol 3Solution Report-Lingol Global optimal sol

Variable X1 X2 X3 value 8.000000 0.000000 0.000000 0.000000 6 . 000000 4.000000 0.000000 0.000000 6 . 000000 Reduced Cost 0.0

New payroll cost = $ 729.6

Increase in payroll cost = 729.6-577.6 = $ 152

d) Complete Linear Program is following:

Let

Let

X1, X2, ... X8 = number of 4-hr shift part-time employees coming to duty at 11 am, 12 pm, ... 6 pm resp.

Y1, Y2, ... Y9 = number of 3-hr shift part-time employees coming to duty at 11 am, 12 pm, ... 7 pm resp.

Objective: Min 7.6*4*(X1+X2+X3+X4+X5+X6+X7+X8)+7.6*3*(Y1+Y2+Y3+Y4+Y5+Y6+Y7+Y8+Y9)

s.t.

X1+Y1 >= 9-1

X1+X2+Y1+Y2 >= 9-1

X1+X2+X3+Y1+Y2+Y3 >= 9-2

X1+X2+X3+X4+Y2+Y3+Y4 >= 3-1

X2+X3+X4+X5+Y3+Y4+Y5 >= 3-2

X3+X4+X5+X6+Y4+Y5+Y6 >= 3-1

X4+X5+X6+X7+Y5+Y6+Y7 >= 6-2

X5+X6+X7+X8+Y6+Y7+Y8 >= 12-2

X6+X7+X8+Y7+Y8+Y9 >= 12-2

X7+X8+Y8+Y9 >= 7-1

X8+Y9 >= 7-1

Xi, Yi >= 0

Solution using LINGO is following:

Lingo 170-Solution Report . Lingo! File Edit Solver Window Help Lindo Model-Lingol 3Solution Report- Lingo1 Global optimal so

Variable X1 Value 1.000000 0.000000 0.000000 0.000000 0.000000 1.000000 0.000000 6.000000 7.000000o 0.000000 0.000000 1.00000

Total payroll cost = $ 494

e) The LP model with the new cost and its solution is following:

Lingo 170-Solution Report . Lingo! File Edit Solver Window Help Lindo Model-Lingol 1G] _İEZ P Solution Report . Lingo! Global

New payroll cost = $ 552.5

The raise in wages increase the cost by = 552.5-494 = $ 58.5